OpenStudy (anonymous):

find the general solution of the differential equation dy/dx=6y^(1/2)

OpenStudy (anonymous):

\[ y' = 6\sqrt{y} \]?

OpenStudy (anonymous):

@wio, I don't think that is correct. the answer to a differential equation's general solution should be in terms of y, not y'

OpenStudy (anonymous):

Okay well it's separable.

OpenStudy (anonymous):

I'm stuck at \[\int\limits_{}^{}\frac{ dy }{ y^{1/2} } = \int\limits_{}^{}6dx\] \[2y ^{1/2} = 6x + c\] I know I need to divide both sides by 2 and then square each side, but does this include the c? (do I do [(6x+c)/2]^2 or just 3x^2 + c?)

OpenStudy (anonymous):

you may assume c/2 = C

OpenStudy (anonymous):

because the constant can be anything

OpenStudy (anonymous):

It's going to be something like \(y=9x^2+6Cx+C^2\)